OpenStudy (anonymous):

Enter the correct solution to the system of equations below as an ordered pair.Do not forget to use the format:"(a,b)". Use the "/" as a fraction bar if necessary. x=-y+3 y=-3x+5

OpenStudy (anonymous):

we have 2 functions: x= function of y f(y) y=function of x f(x) we know that the solution of system will make both of these true so at the solution, =we can actually combine the 2 functions: x= -y+3 y=-3x+5 by substituing y into f(x) we get x=-(-3x+5)+3 we can simplify this to get: x=3x -2 we put all of the "x"s on one side: -2x=-2 x=1 so now we have x, but we still need y we can get this using our equations: x=-y+3 1-3= -y -2=-y y=2 so the solution of out system is going to be (x,y) = (1,2)

OpenStudy (anonymous):

Thanks=)